# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v2

@@ A.3.7 The max function is not Lipschitz continuous with respect to EMD
$f_{ij}=\begin{cases}1&\text{if }i=j\\ 0&\text{if }i\neq j\\ \end{cases}$
Therefore, the EMD between $X$ and $Y$ is equal to $1$. Then, we have that
$||\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}||$ $=||\mathbf{v}_{m}-\mathbf{u}_{m}||$
$=m\cdot 1$
$=m\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}||\mathbf{v}_{i}-\mathbf{u}_{j}||$
$>L\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}]_{ij}||\mathbf{v}_{i}-\mathbf{u}_{j}||$
@@
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $||\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}||>L\,d_{\text{EMD}}(X,Y)$, which is a contradiction. Therefore, the max function is not Lipschitz continuous with respect to EMD.
A.3.8 The max function is not Lipschitz continuous with respect to the matching distance
Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we have $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=||\mathbf{v}_{2}||=\sqrt{d}\epsilon$. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We also have that:
$\Big{|}\Big{|}f_{\text{max}}(X)-f_{\text{max}}(Y)\Big{|}\Big{|}$ $=||\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}||$
$=||\mathbf{v}_{\text{max}}-\mathbf{u}_{1}||$
$=\Big{|}\Big{|}\big{(}\max(-c,\epsilon),\max(-c,\epsilon),\ldots,\max(-c,\epsilon)\big{)}^{\top}-(-c,-c,\ldots,-c)^{\top}\Big{|}\Big{|}$
$=||(\epsilon,\epsilon,\ldots,\epsilon)^{\top}-(-c,-c,\ldots,-c)^{\top}||$
@@ A.3.8 The max function is not Lipschitz continuous with respect to the matching distance
For any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $||f_{\text{max}}(X)-f_{\text{max}}(Y)||>L\,d_{M}(X,Y)$. Based on the above inequality, the max function is not Lipschitz continuous with respect to the matching distance.
A.3.9 The max function is Lipschitz continuous with respect to the Hausdorff distance
Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ denote two multisets of vectors. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We will show that $\forall k\in[d]$, we have that $\big{|}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{|}\leq d_{H}(X,Y)$. By contradiction, we assume that that there is some $k\in[d]$ such that $\big{|}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{|}>d_{H}(X,Y)$. Without loss of generality, we also assume that $[\mathbf{v}_{\text{max}}]_{k}\geq[\mathbf{u}_{\text{max}}]_{k}$, and therefore $\big{|}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{|}=[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}$.
Since $[\mathbf{u}_{\text{max}}]_{k}\geq[\mathbf{u}_{j}]_{k}$, $\forall j\in[n]$, we have that $[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\leq[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{j}]_{k}$, $\forall j\in[n]$. Since by our assumption above, $\big{|}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{|}=[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}>d_{H}(X,Y)$, it follows that
$[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y),\quad\forall j\in[n]$ (1)
Note that there is at least one vector $\mathbf{v}_{i}\in X$ such that $[\mathbf{v}_{i}]_{k}=[\mathbf{v}_{\text{max}}]_{k}$. From equation (1), we have for this vector that $[\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y)$, $\forall j\in[n]$. Then, we have $||\mathbf{v}_{i}-\mathbf{u}_{j}||=\sqrt{\big{(}[\mathbf{v}_{i}]_{1}-[\mathbf{u}_{j}]_{1}\big{)}^{2}+\ldots+\big{(}[\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}\big{)}^{2}+\ldots+\big{(}[\mathbf{v}_{i}]_{d}-[\mathbf{u}_{j}]_{d}\big{)}^{2}}\geq\sqrt{\big{(}[\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}\big{)}^{2}}=[\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y)$, $\forall j\in[n]$. We thus have that $\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{u}_{j}||>d_{H}(X,Y)$ which is a contradiction since $\min_{j\in[n]}||\mathbf{v}_{i}-\mathbf{u}_{j}||\leq\max_{\imath\in[m]}\min_{j\in[n]}||\mathbf{v}_{\imath}-\mathbf{u}_{j}||=h(X,Y)\leq d_{H}(X,Y)$. Therefore, we have that $\big{|}[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big{|}\leq d_{H}(X,Y)$.
@@ A.3.9 The max function is Lipschitz continuous with respect to the Hausdorff distance
$\Big{|}\Big{|}f_{\text{max}}(X)-f_{\text{max}}(Y)\Big{|}\Big{|}$ $=||\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}||$
@@
A.4 Proof of Lemma 3.2
A.4.1 The mean function is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $X\in\mathcal{S}(\mathbb{R}^{d}\setminus\{\mathbf{0}\})$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. By Proposition 2.3, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$. By Theorem 3.1, we have that:
$\Big{|}\Big{|}f_{\text{mean}}(X)-f_{\text{mean}}(Y)\Big{|}\Big{|}$ $\leq d_{\text{EMD}}(X,Y)$
@@ A.4.1 The mean function is Lipschitz continuous with respect to the matching distance
A.4.2 The sum function is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. By Proposition 2.3, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$. By Theorem 3.1, we have that:
$\Big{|}\Big{|}f_{\text{sum}}(X)-f_{\text{sum}}(Y)\Big{|}\Big{|}$ $\leq d_{M}(X,Y)$
@@ A.4.2 The sum function is Lipschitz continuous with respect to EMD
A.4.3 The max function is Lipschitz continuous with respect to the matching distance
@@
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$, $\forall X\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}\}\!\!\}\in\mathcal{X}$, and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{M}\}\!\!\}\in\mathcal{X}$ denote two multisets. Then, we have that:
$\Big{|}\Big{|}f_{\text{max}}(X)-f_{\text{max}}(Y)\Big{|}\Big{|}=\sqrt{\big{(}[\mathbf{v}_{\text{max}}]_{1}-[\mathbf{u}_{\text{max}}]_{1}\big{)}^{2}+\big{(}[\mathbf{v}_{\text{max}}]_{2}-[\mathbf{u}_{\text{max}}]_{2}\big{)}^{2}+\ldots+\big{(}[\mathbf{v}_{\text{max}}]_{d}-[\mathbf{u}_{\text{max}}]_{d}\big{)}^{2}}$ (2)

### v3

@@ B.3.7 The max function is not Lipschitz continuous with respect to EMD
[…]
Therefore, the EMD between $X$ and $Y$ is equal to $1/m$. Then, we have that
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|$ $=\|\mathbf{v}_{m}-\mathbf{u}_{m}\|$
$=m\cdot\frac{1}{m}$
$=m\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}^{*}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
$>L\sum_{i=1}^{m}\sum_{j=1}^{m}[\mathbf{F}^{*}]_{ij}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|$
@@
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\|>L\,d_{\text{EMD}}(X,Y)$, which is a contradiction. Therefore, the max function is not Lipschitz continuous with respect to EMD.
B.3.8 The max function is not Lipschitz continuous with respect to the matching distance
Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We also have that:
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|$ $=\|\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}\|$
$=\|\mathbf{v}_{\text{max}}-\mathbf{u}_{1}\|$
$=\Big\|\big(\max(-c,\epsilon),\max(-c,\epsilon),\ldots,\max(-c,\epsilon)\big)^{\top}-(-c,-c,\ldots,-c)^{\top}\Big\|$
$=\|(\epsilon,\epsilon,\ldots,\epsilon)^{\top}-(-c,-c,\ldots,-c)^{\top}\|$
@@ B.3.8 The max function is not Lipschitz continuous with respect to the matching distance
For any $L>0$, there exist $X,Y\in\mathcal{S}(\mathbb{R}^{d})$ such that $\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\|>L\,d_{M}(X,Y)$. Based on the above inequality, the max function is not Lipschitz continuous with respect to the matching distance.
B.3.9 The max function is Lipschitz continuous with respect to the Hausdorff distance
Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}\!\!\}$ and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{n}\}\!\!\}$ denote two multisets of vectors. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We will show that $\forall k\in[d]$, we have that $\big|[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big|\leq d_{H}(X,Y)$. By contradiction, we assume that that there is some $k\in[d]$ such that $\big|[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big|>d_{H}(X,Y)$. Without loss of generality, we also assume that $\mathbf{v}_{\text{max}}]_{k}\geq[\mathbf{u}_{\text{max}}]_{k}$, and therefore $\big|[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big|=[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}$.
Since $\mathbf{u}_{\text{max}}]_{k}\geq[\mathbf{u}_{j}]_{k}$, $\forall j\in[n]$, we have that $\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\leq[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{j}]_{k}$, $\forall j\in[n]$. Since by our assumption above, $\big|[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big|=[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}>d_{H}(X,Y)$, it follows that
$\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y),\quad\forall j\in[n]$ (1)
Note that there is at least one vector $\mathbf{v}_{i}\in X$ such that $\mathbf{v}_{i}]_{k}=[\mathbf{v}_{\text{max}}]_{k}$. From equation 1, we have for this vector that $\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y)$, $\forall j\in[n]$. Then, we have $\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\sqrt{\big([\mathbf{v}_{i}]_{1}-[\mathbf{u}_{j}]_{1}\big)^{2}+\ldots+\big([\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}\big)^{2}+\ldots+\big([\mathbf{v}_{i}]_{d}-[\mathbf{u}_{j}]_{d}\big)^{2}}\geq\sqrt{\big([\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}\big)^{2}}=[\mathbf{v}_{i}]_{k}-[\mathbf{u}_{j}]_{k}>d_{H}(X,Y)$, $\forall j\in[n]$. We thus have that $\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|>d_{H}(X,Y)$ which is a contradiction since $\min_{j\in[n]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|\leq\max_{\imath\in[m]}\min_{j\in[n]}\|\mathbf{v}_{\imath}-\mathbf{u}_{j}\|=h(X,Y)\leq d_{H}(X,Y)$. Therefore, we have that $\big|[\mathbf{v}_{\text{max}}]_{k}-[\mathbf{u}_{\text{max}}]_{k}\big|\leq d_{H}(X,Y)$.
@@ B.3.9 The max function is Lipschitz continuous with respect to the Hausdorff distance
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|$ $=\|\mathbf{v}_{\text{max}}-\mathbf{u}_{\text{max}}\|$
@@
B.4 Proof of Lemma 3.2
B.4.1 The mean function is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. By Proposition 2.3, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$. By Theorem 3.1, we have that:
$\Big\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\Big\|$ $\leq d_{\text{EMD}}(X,Y)$
@@ B.4.1 The mean function is Lipschitz continuous with respect to the matching distance
B.4.2 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the mean function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of $3$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
$\Big\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\Big\|$ $=\bigg\|\frac{1}{3}\sum_{i=1}^{3}\mathbf{v}_{i}-\frac{1}{3}\sum_{i=1}^{3}\mathbf{u}_{1}\bigg\|$
$=\bigg\|\frac{1}{3}(\mathbf{v}_{1}-\mathbf{u}_{1}+\mathbf{v}_{2}-\mathbf{u}_{2}+\mathbf{v}_{3}-\mathbf{u}_{3})\bigg\|$
$=\frac{1}{3}\|\mathbf{v}_{3}-\mathbf{u}_{3}\|$
$=\frac{1}{3}\Big\|\Big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\Big)^{\top}-\Big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\Big)^{\top}\Big\|$
$=\frac{1}{3}\|(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}\|$
$=\frac{1}{3}\sqrt{d(c+\epsilon)^{2}}$
$=\frac{1}{3}\sqrt{d}(c+\epsilon)$
$>\frac{1}{3}\sqrt{d}(3L\epsilon+\epsilon)$
$>L\sqrt{d}\epsilon$
$>L\frac{\sqrt{d}\epsilon}{2}$
$=L\,d_{H}(X,Y)$
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{X}$ such that $\|f_{\textsc{mean}}(X)-f_{\textsc{mean}}(Y)\|>L\,d_{H}(X,Y)$, which is a contradiction. Therefore, the mean function is not Lipschitz continuous with respect to the Hausdorff distance even when it is restricted to inputs from set $\mathcal{X}$.
B.4.3 The sum function is Lipschitz continuous with respect to EMD
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X,Y\in\mathcal{X}$ denote two multisets. By Proposition 2.3, we have that $d_{M}(X,Y)=Md_{\text{EMD}}(X,Y)$. By Theorem 3.1, we have that:
$\Big\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\Big\|$ $\leq d_{M}(X,Y)$
@@ B.4.3 The sum function is Lipschitz continuous with respect to EMD
B.4.4 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
@@
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the sum function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of $3$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
$\Big\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\Big\|$ $=\bigg\|\sum_{i=1}^{3}\mathbf{v}_{i}-\sum_{i=1}^{3}\mathbf{u}_{1}\bigg\|$
$=\|\mathbf{v}_{1}-\mathbf{u}_{1}+\mathbf{v}_{2}-\mathbf{u}_{2}+\mathbf{v}_{3}-\mathbf{u}_{3}\|$
$=\|\mathbf{v}_{3}-\mathbf{u}_{3}\|$
$=\Big\|\Big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\Big)^{\top}-\Big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\Big)^{\top}\Big\|$
$=\|(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}\|$
$=\sqrt{(c+\epsilon)^{2}+(c+\epsilon)^{2}+\ldots+(c+\epsilon)^{2}}$
$=\sqrt{d(c+\epsilon)^{2}}$
$=\sqrt{d}(c+\epsilon)$
$>\sqrt{d}(L\epsilon+\epsilon)$
$>L\frac{\sqrt{d}\epsilon}{2}$
$=L\,d_{H}(X,Y)$
Therefore, for any $L>0$, there exist $X,Y\in\mathcal{X}$ such that $\|f_{\textsc{sum}}(X)-f_{\textsc{sum}}(Y)\|>L\,d_{H}(X,Y)$, which is a contradiction. Therefore, the sum function is not Lipschitz continuous with respect to the Hausdorff distance even when it is restricted to inputs from set $\mathcal{X}$.
B.4.5 The max function is Lipschitz continuous with respect to the matching distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Let $X=\{\!\!\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}\}\!\!\}\in\mathcal{X}$, and $Y=\{\!\!\{\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{M}\}\!\!\}\in\mathcal{X}$ denote two multisets. Then, we have that:
$\Big\|f_{\textsc{max}}(X)-f_{\textsc{max}}(Y)\Big\|=\sqrt{\big([\mathbf{v}_{\text{max}}]_{1}-[\mathbf{u}_{\text{max}}]_{1}\big)^{2}+\big([\mathbf{v}_{\text{max}}]_{2}-[\mathbf{u}_{\text{max}}]_{2}\big)^{2}+\ldots+\big([\mathbf{v}_{\text{max}}]_{d}-[\mathbf{u}_{\text{max}}]_{d}\big)^{2}}$ (2)